Maximum Accounting firm currently has two offices in Southern California. The following table shows each office’s account numbers and profits. The firm is planning to open a new office in Northridge. The expected number of accounts for the new Northridge office is 3. Please use linear regression method to forecast the profits that would be generated from the new Northridge office.

Office Pasadena Long Beach Northridge
# of accounts (x) 8 10 3
Profit (y), in millions $ 14 20 ???

Inputs will be exact numbers.

  1. What is the parameter value of b in the linear regression function?
  2. Assume the parameter in the linear regression function b = - 0.5. Please use the assumed b value and other information in the problem to forecast the profits from Northridge office.
Answer #1

Note: In case of any doubt, please mention it in the comments. Thank you

. X Y XY х2 8 10 14 20 119 200 64 100 EX = 18; & Y= 34; £XY=312; 2X=164 Sloppe (b) = n(Exy) - (x) (Y) m (4x2) -(x)? =( 9 x 31is: Linear regression equation yza a tbx y = 21.5-0.50 Given a=3 So, 21.5- (0.5x3) Therefore, forecast of profits = 20 Scanne